A plane flies 720 mi against a steady 30 mi.h headwind and then returns to the same point with the wind. If the entire trip takes 10 hr what is the planes speed in still air?: A plane flies 720 mi against a steady 30 mi.h headwind and then returns to the same point with the wind. If the entire trip takes 10 hr what is the planes speed in still air?

Answer by ptaylor(1894)


Let x=speed (rate) of the plane in still air
speed against headwind =x-30
speed with headwind=x+30
Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Time required travelling against the headwind=720/(x-30)
Time require travelling with the headwind=720/(x+30)
Now we are told that the total time was 10 hours, So:
720/(x-30)+720/(x+30)=10 multiply both sides by (x-30)(x+30)
720(x+30)+720(x-30)=10(x-30)(x+30) divide both sides by 10
72(x+30)+72(x-30)=(x-30)(x+30) get rid of parens
72x+2160+72x-2160=x^2-900 collect like terms
144x=x^2-900 subtract 144x from both sides
144x-144x=x^2-144x-900 simplify
x^2-144x-900=0 quadratic in standard form and this can be factored
(x-150)(x+6)=0
x=150mph-------------------------------------rate in still air
x=-6----------------discount negative value for speed
CK
720/(150-30)+720/(150+30)=10
6+4=10
10=10

Hope this helps---ptaylor